### v1

@@ 2.5.2 Case 2: Training with acceleration measurements only
$J(\theta)=\frac{1}{N}\left\|\dot{x}^{p}-x_{t}^{p}\right\|_{2}^{2}+\frac{1}{N}\left\|\dot{x}_{t}^{p}-\ddot{x}^{m}\right\|_{2}^{2}+\frac{1}{N}\left\|\dot{x}_{t}^{p}+g^{p}+\Gamma\ddot{x}_{g}\right\|_{2}^{2},$
where only acceleration measurements $\ddot{x}^{m}$ contribute to the loss due to data availability. The model is trained on 50 randomly selected datasets and tested on 50 unseen datasets.
Figure 8 illustrates the displacement time histories predicted by the proposed PhyULSTM model alongside those from the baseline PhyCNN and the ground truth numerical solutions for three distinct, previously unseen seismic excitation inputs. The PhyULSTM predictions (left) closely track the true system response (blue), accurately capturing both the amplitude and phase of complex oscillatory behavior inherent to nonlinear structural dynamics. In contrast, the PhyCNN predictions (right) exhibit noticeable deviations, particularly during peak displacement intervals and rapid transient phases, indicating an underestimation of key dynamic features. Figure 9 further highlights the predictive fidelity of PhyULSTM by comparing the velocity ($\dot{x}$) and nonlinear restoring force ($g$) time histories for a representative unseen seismic event. The model faithfully reproduces the intricate temporal variations and sharp fluctuations characteristic of nonlinear hysteretic behavior, which are only partially captured by PhyCNN.
Figure 9: PhyULSTM accurately reproduces velocity ($\dot{x}$) and restoring force ($g$) time histories, closely matching the ground truth obtained from numerical integration. The left column shows the responses predicted by PhyULSTM, and the right column shows those predicted by PhyCNN. Subplots are labeled 4a, 4b, 4c, and 4d for reference: 4a and 4c correspond to velocity and normalized restoring force predicted by PhyULSTM, while 4b and 4d correspond to the same predicted by PhyCNN. The model effectively captures the nonlinear system response including sharp fluctuations, which alternative methods fail to capture effectively.
@@
2.6 Experimental Validation
@@
To validate this data-driven variant of the PhyULSTM model, we utilize recorded seismic response data from a six-story reinforced concrete hotel building located in San Bernardino, California. The structure, designed in 1970, has been instrumented with nine accelerometers on the first, third, and roof levels in both principal directions. The dataset, curated by the Center for Engineering Strong Motion Data (CESMD) [haddadi2012report], comprises earthquake recordings collected between 1987 and 2018. These events span a broad range of magnitudes and frequency content, thereby providing a representative and challenging benchmark for data-driven structural response modeling. For consistency and reproducibility, the dataset partitioning follows the protocol established in Zhang et al. [phycnn], which involves stratified sampling into training, validation, and test subsets. The training objective for the data-driven PhyULSTM model is to learn a mapping from ground acceleration inputs to structural displacement outputs.
Figure 11: Predicted displacement response at the 3rd floor (DOF-1) of the six-story instrumented hotel building in San Bernardino, California, subjected to the 2014 Big Bear Lake and 2016 Loma Linda earthquakes. The model was trained using historical seismic data from 1987–2018, relying solely on acceleration measurements from strategically placed sensors, without incorporating any explicit physical constraints in the loss function. Despite this, the PhyULSTM model successfully captures the dynamic behavior of the structure, including transient displacement patterns and resonant oscillations, aligning well with sensor-derived ground truth.
@@
$J(\theta)=\frac{1}{N}\left\|\ddot{x}^{m}-x_{tt}^{p}\right\|_{2}^{2}$
@@ 3 Conclusion
This study introduces a hybrid deep learning framework, PhyULSTM, that integrates the architectural strengths of U-net and LSTM within a physics-informed training regime for the surrogate modeling of nonlinear structural systems under seismic excitation. When available, governing physical laws are incorporated directly into the loss function to guide learning, mitigate overfitting, and enhance generalization, particularly in data-scarce environments. Moreover, the model outperforms existing seismic response prediction models like PhyCNN and is also capable of operating without explicit physical constraints, enabling its application to systems with incomplete or unknown constitutive characteristics. Extensive numerical experiments on nonlinear single-degree-of-freedom systems and experimental validation using sensor data from a mid-rise concrete building in San Bernardino demonstrate that the PhyULSTM framework outperforms existing physics-informed CNN architectures. It exhibits superior predictive accuracy, higher correlation in regression analysis, and robust performance under varied excitation characteristics, including amplitude and frequency content. These results underscore the framework’s ability to generalize structural response behavior under seismic loading, even when trained on limited and noisy data.
The modular nature of the PhyULSTM model allows it to be extended to other structural typologies, including bridges, tunnels, and other civil infrastructure. Furthermore, the combination of U-net and LSTM makes the architecture well-suited for broader time-series modeling tasks beyond structural engineering, such as applications in signal processing and sequential decision-making. Overall, this study establishes PhyULSTM as a reliable, interpretable, and physically grounded metamodeling tool that advances the state of surrogate modeling for structural dynamics and provides a foundation for future developments in hybrid physics-informed machine learning.

### v2

@@ 3.2 Case 2: Training with acceleration measurements only
$J(\theta)=J_{D}(\theta)+J_{P}(\theta)$
where
$J_{P}(\theta)=\frac{1}{N}\sum_{i=1}^{N}\left\|\dot{x}^{p(i)}-x_{t}^{p(i)}\right\|_{2}^{2}+\frac{1}{N}\sum_{i=1}^{N}\left\|\dot{x}_{t}^{p(i)}+g^{p(i)}+\Gamma\ddot{x}^{(i)}_{g}\right\|_{2}^{2}$
and
$J_{D}(\theta)=\frac{1}{N}\sum_{i=1}^{N}\left\|\dot{x}_{t}^{p(i)}-\ddot{x}^{m(i)}\right\|_{2}^{2}$
Only acceleration measurements $\ddot{x}^{m}$ contribute to the loss due to data availability. The model is trained on 50 randomly selected datasets and tested on 50 unseen datasets.
Figure 10 illustrates the displacement time histories predicted by the proposed PhyULSTM model alongside those from the baseline PhyCNN and the ground truth numerical solutions for three distinct, previously unseen seismic excitation inputs. The PhyULSTM predictions (left) closely track the true system response (blue), accurately capturing both the amplitude and phase of complex oscillatory behavior inherent to nonlinear structural dynamics. In contrast, the PhyCNN predictions (right) exhibit noticeable deviations, particularly during peak displacement intervals and rapid transient phases, indicating an underestimation of key dynamic features. Figure 11 further highlights the predictive fidelity of PhyULSTM by comparing the velocity ($\dot{x}$) and acceleration ($\ddot{x}$) time histories for a representative unseen seismic excitation. The close agreement between the predicted and reference responses demonstrates the model’s capability to accurately capture the underlying nonlinear dynamic behavior of the system.
Figure 11: PhyULSTM accurately reproduces velocity ($\dot{x}$) and acceleration ($\ddot{x}$) time histories, closely matching the ground truth obtained from numerical integration. The left column shows the responses predicted by PhyULSTM, and the right column shows those predicted by PhyCNN. Subplots are labeled 4a, 4b, 4c, and 4d for reference: 4a and 4c correspond to velocity and acceleration predicted by PhyULSTM, while 4b and 4d correspond to the same quantities predicted by PhyCNN.
@@
3.2.1 Global error comparison
Table 3 summarizes the prediction errors obtained when only acceleration measurements are available during training. The PhyULSTM model achieves a mean RMSE of $1.02\times 10^{-2}$ compared to $2.23\times 10^{-2}$ for PhyCNN, corresponding to an error reduction of approximately $54\%$. The improvement is even more pronounced in the relative $L_{2}$ error, where PhyULSTM achieves $20.95\%$ compared to $51.50\%$ for PhyCNN. The median errors follow the same trend, confirming the robustness of the improvement across all test cases.
Table 3: Global error metrics for Case 2 (acceleration-only training).
Model Mean RMSE Median RMSE Mean Rel-$L_{2}$ Median Rel-$L_{2}$
PhyULSTM $1.0195\times 10^{-2}$ $6.7850\times 10^{-3}$ $20.95\%$ $14.90\%$
PhyCNN $2.2338\times 10^{-2}$ $1.6906\times 10^{-2}$ $51.50\%$ $34.04\%$
3.2.2 Frequency-domain comparison
The Fourier amplitude spectra of the predicted responses are shown in Fig. 13. The spectra confirm that the structural response remains dominated by low-frequency components (<2<2 Hz). Both models capture the dominant frequency locations and general spectral decay.
Figure 13: Comparison of Fourier amplitude spectra of reference and predicted responses for multiple test samples under acceleration-only training.
However, noticeable differences appear in the peak spectral amplitudes. In several cases, the PhyCNN model either overestimates or underestimates the peak amplitude and shows mild spectral spreading at higher frequencies. In contrast, the PhyULSTM predictions consistently show closer agreement with the reference spectra and preserve the dominant peak magnitudes more accurately. This behavior indicates improved reconstruction of structural dynamics when temporal dependencies are modeled.
3.2.3 Peak response error analysis
The distribution of peak response errors is presented in Fig. 14. The PhyULSTM model achieves a mean peak error of $4.37\%$ and a median error of $2.17\%$, while PhyCNN produces significantly larger values of $12.83\%$ and $11.88\%$, respectively.
Figure 14: Distribution of peak response errors for acceleration-only training.
In addition to the lower central tendency, the PhyULSTM predictions exhibit a much tighter error distribution and substantially fewer extreme outliers.
4 Application in Response Prediction of a Bouc–Wen Model with Rate-Dependent Hysteresis
The proposed PhyULSTM framework is applied to the response prediction of a nonlinear single-degree-of-freedom (SDOF) system governed by a rate-dependent Bouc–Wen hysteresis model [54]. This benchmark problem is widely used to assess the capability of data-driven and physics-informed models in capturing nonlinear, history-dependent structural behavior.
The equation of motion of the SDOF system subjected to ground excitation is expressed as
$m\,\ddot{u}(t)+c\,\dot{u}(t)+\lambda k\,u(t)+(1-\lambda)k\,r(t)=-m\,a_{g}(t)$ (19)
where $u(t)$ denotes the displacement response, $\dot{u}(t)$ and $\ddot{u}(t)$ are the velocity and acceleration, respectively, $m$, $c$, and $k$ are the mass, damping, and stiffness coefficients, $r(t)$ is the hysteretic internal variable, $\lambda\in(0,1]$ is the post-yield stiffness ratio, and $a_{g}(t)$ is the ground acceleration.
Normalizing Eq. (19) by the mass $m$ yields
$\ddot{u}(t)+2\zeta\omega_{n}\,\dot{u}(t)+\lambda\omega_{n}^{2}\,u(t)+(1-\lambda)\omega_{n}^{2}\,r(t)=-a_{g}(t)$ (20)
where $\omega_{n}=\sqrt{k/m}$ is the natural frequency and $\zeta=c/(2\sqrt{km})$ is the damping ratio. The corresponding mass-normalized restoring force is defined as
$g(t)=2\zeta\omega_{n}\,\dot{u}(t)+\lambda\omega_{n}^{2}\,u(t)+(1-\lambda)\omega_{n}^{2}\,r(t),$ (21)
which leads to the compact equation of motion
$\ddot{u}(t)+g(t)=-a_{g}(t).$ (22)
The evolution of the hysteretic variable $r(t)$ follows the rate-dependent Bouc–Wen differential equation
$\dot{r}(t)=\dot{u}(t)-\alpha|\dot{u}(t)||r(t)|^{n-1}r(t)-\beta\dot{u}(t)|r(t)|^{n}$ (23)
where $\alpha$ controls the size and shape of the hysteresis loop, $\beta$ governs the smoothness of the transition between elastic and inelastic regimes, and $n\geq 1$ determines the sharpness of yielding. The initial condition is set as $r(0)=0$.
Table 4: Bouc–Wen System Parameters
Parameter Symbol Value
Mass $m$ 500 kg
Damping coefficient $c$ 0.35 kN·s/m
Stiffness $k$ 25 kN/m
Bouc–Wen parameter $\alpha$ 2.0
Bouc–Wen parameter $\beta$ 2.0
Bouc–Wen exponent $n$ 3.0
Post-yield stiffness ratio $\lambda$ 0.5
System Parameters
The Bouc–Wen system parameters adopted in this study are summarized in Table 4. The corresponding derived system properties are the natural frequency $\omega_{n}=7.0711$ rad/s, natural period $T_{n}=0.8886$ s, damping ratio $\zeta=0.007$, and an approximate yield displacement of $u_{y}\approx 0.01$ m.
Loss Function Formulation
In this case, the PhyULSTM architecture directly predicts velocity and mass-normalized restoring force, while displacement is obtained via numerical time integration and acceleration is computed using finite-difference differentiation denoted by ${\int}\dot{u}$ and $\dot{u}_{t}$ respectively. Model training is guided by a composite loss function that combines data fidelity with physics-based constraints:
$\mathcal{L}_{\text{total}}=\mathcal{L}_{\text{data}}+\mathcal{L}_{\text{physics}}$ (24)
where each term contributes to a distinct aspect of model supervision.
The data loss penalizes discrepancies between predicted and reference structural responses across displacement, velocity, and acceleration. For a batch of $N$ ground motion samples, the data loss is defined as
$\mathcal{L}_{\text{data}}=\frac{1}{N}\sum_{i=1}^{N}\left\|\int\dot{u}^{p(i)}-u^{true(i)}\right\|_{2}^{2}+\frac{1}{N}\sum_{i=1}^{N}\left\|\dot{u}^{p(i)}-\dot{u}^{true(i)}\right\|_{2}^{2}+\frac{1}{N}\sum_{i=1}^{N}\left\|\dot{u}_{t}^{p(i)}-\ddot{u}^{true(i)}\right\|_{2}^{2}$ (25)
The physics loss enforces dynamic equilibrium. Here, the superscript $p$ denotes the predicted response. Using $\Phi$ to denote the finite-difference operator, the equilibrium residual for the $i$-th sample is defined as
$r^{(i)}=\dot{u}_{t}^{p(i)}+g^{(i)}_{\text{pred}}+a_{g}^{(i)}$
with $\dot{u}_{t}=\Phi\dot{u}$. The physics loss is given by
$\mathcal{L}_{\text{physics}}=\frac{1}{N_{c}}\sum_{i=1}^{N_{c}}\|r^{(i)}\|_{2}^{2}$ (26)
where $N$ is the number of training samples and $N_{c}$ denotes the number of collocation points. Optional weighting parameters may be introduced to balance the relative importance of data fidelity and physics enforcement.Here, the superscript $p$ denotes the predicted response.
If desired, the different components may be balanced using scalar weights $\lambda_{\text{data}}$, $\lambda_{\text{phys}}$. The weighted total loss reads:
$\mathcal{L}_{\text{total}}\;=\;\lambda_{\text{data}}\mathcal{L}_{\text{data}}\;+\;\lambda_{\text{phys}}\mathcal{L}_{\text{physics}}$ (27)
Choosing the weights allows control over the relative importance of data fidelity, and physics enforcement during training.
The system is discretized using a time step $\Delta t=0.02$ s (sampling frequency $50$ Hz). The Bouc–Wen hysteresis equation (23) is integrated using a forward Euler scheme. A synthetic dataset consisting of 100 independent seismic response simulations is generated by exciting the SDOF system with band-limited white noise (BLWN) ground motions of varying intensities. Each simulation spans 30 s, resulting in 1501 time steps per record. Ten BLWN input–response pairs are randomly selected as “known” datasets for training and validation using an 80/20 split, while the remaining samples are treated as “unknown” datasets for evaluating generalization. In addition, 50 collocation samples containing only ground motion records are included during training to strengthen physics enforcement. The dataset is publicly available at https://github.com/zhry10/PhyLSTM.
Figure 15: Predicted vs reference responses — displacement, velocity, acceleration, and normalized restoring force for an unseen earthquake from the test dataset.
Unlike the PhyLSTM2 and PhyLSTM3 frameworks [54], which employ multiple LSTM networks to separately model state variables and hysteresis evolution, the proposed PhyULSTM uses a single deep LSTM coupled with a 1D U-Net feature extractor. The network is trained for 5,000 epochs using the Adam optimizer with a learning rate of $10^{-3}$, followed by an additional 5,000 epochs with a reduced learning rate of $10^{-4}$. This unified architecture enforces internal consistency among state variables, mitigates drift accumulation, and improves computational efficiency.
Figure 15 presents displacement, velocity, acceleration, and normalized restoring force responses for an unseen earthquake record. The PhyULSTM predictions closely match the numerical reference across all response quantities. Quantitative performance comparisons further highlight the superiority of PhyULSTM. For displacement prediction, correlation coefficients range from $0.911$ to $1.000$, compared to 0.770.770.990.99 for PhyLSTM3 and 0.190.190.850.85 for PhyLSTM2. For velocity, acceleration, and restoring force, PhyULSTM consistently achieves correlation coefficients exceeding $0.99$. The probability distributions of correlation coefficients for unseen earthquakes are shown in Figure 16.
A key distinction between PhyULSTM and PhyLSTM3 lies in the treatment of hysteresis. While PhyLSTM3 explicitly models the hysteretic variable $r(t)$ using a dedicated LSTM network, PhyULSTM learns hysteretic behavior implicitly through its hierarchical feature extraction and temporal memory. These results indicate that, when appropriately regularized by physics-based constraints, deep neural networks can learn physically meaningful, history-dependent representations without explicitly encoding the governing constitutive equations.
Figure 16: Correlation coefficient versus probability distributions for the predicted displacement (a), velocity (b), acceleration (c), and normalised restoring force (d) responses corresponding to unseen earthquakes from the test dataset.
5 Application in response prediction of a six-story reinforced concrete building
@@
To validate this data-driven variant of the PhyULSTM model, we utilize recorded seismic response data from a six-story reinforced concrete hotel building located in San Bernardino, California. The structure, designed in 1970, has been instrumented with nine accelerometers on the first, third, and roof levels in both principal directions. The dataset, curated by the Center for Engineering Strong Motion Data (CESMD) [12], comprises earthquake recordings collected between 1987 and 2018. These events span a broad range of magnitudes and frequency content, thereby providing a representative and challenging benchmark for data-driven structural response modeling. For consistency and reproducibility, the dataset partitioning follows the protocol established in Zhang et al. [53], which involves stratified sampling into training, validation, and test subsets. The training objective for the data-driven PhyULSTM model is to learn a mapping from ground acceleration inputs to structural displacement outputs.
Figure 17: Predicted displacement response at the 3rd floor (DOF-1) of the six-story hotel building in San Bernardino, California, subjected to the 2014 Big Bear Lake and 2016 Loma Linda earthquakes.
@@
$J_{D}(\theta)=\frac{1}{N}\sum_{i=1}^{N}\left\|\ddot{x}^{m(i)}-x_{tt}^{p(i)}\right\|_{2}^{2}$
@@
5.1 Error Distribution Comparison
To evaluate the prediction accuracy of the proposed framework, the probability density function (PDF) of the normalised displacement error is analysed and compared with the results reported for the PhyCNN model. The normalised error is defined as the difference between the predicted and reference responses divided by the maximum absolute reference displacement.
$e=\frac{y_{\text{true}}-y_{\text{pred}}}{\max(|y_{\text{true}}|)},$ (28)
where $y_{\text{true}}$ and $y_{\text{pred}}$ denote the measured and predicted displacements, respectively. The probability density function (PDF) of this normalized error is estimated via kernel density estimation and plotted for both the 3rd floor and roof levels in Figure19.
5.1.1 Confidence Interval Comparison
Comparison of prediction error confidence intervals within $\pm 5\%$ is presented in Table 5. For the PhyCNN model, the error distribution shows that the prediction error is predominantly concentrated within $\pm 5\%$, with confidence intervals (CI) of 97% for the 3rd floor and 93% for the roof. The PDFs are sharply peaked near zero with minimal tails beyond the $\pm 5\%$ bounds, demonstrating high point-wise accuracy across the prediction set. In comparison, the proposed PhyULSTM achieves comparable or slightly superior performance on the same benchmark case (six-story hotel building, acceleration-only inputs, and identical data partitioning via K-means clustering). Specifically, the CI within the $\pm 5\%$ error band is 99.89% for the 3rd floor and 98.05% for the roof (exceeding PhyCNN’s 93%). The error PDFs for both levels remain tightly concentrated around zero, with lower dispersion at the roof level, as evidenced by the higher CI and narrower peak in the plot. This indicates excellent agreement with the reference responses, particularly for the upper stories where long-range temporal dependencies are more pronounced.
Model 3rd Floor Roof
PhyCNN 97% 93%
PhyULSTM 99.89% 98.05%
Table 5: Comparison of prediction error confidence intervals within $\pm 5\%$.
Figure 19: Probability density function of normalised prediction error for the 3rd floor and roof displacement responses. The dashed vertical lines indicate the $\pm 5\%$ error bounds. The proposed model achieves confidence intervals of 99.89% and 98.05% for the 3rd floor and roof, respectively.
5.1.2 Global error comparison
Both models achieve very small absolute errors, indicating that the overall structural response is well captured.
PhyULSTM achieves a mean RMSE of $1.13\times 10^{-3}$ compared to $1.44\times 10^{-3}$ for PhyCNN, corresponding to an improvement of approximately $22\%$. The relative $L_{2}$ error is also slightly lower for PhyULSTM ($0.23\%$ vs $0.28\%$), confirming its superior accuracy. Although the difference in global errors is modest due to the high accuracy of both models, the advantage of PhyULSTM becomes more evident when peak responses are examined.
The global error metrics are summarized in Table  6.
Model Mean RMSE Median RMSE Mean Rel-$L_{2}$ Median Rel-$L_{2}$
PhyULSTM $1.1289\times 10^{-3}$ $1.0152\times 10^{-3}$ 0.23% 0.26%
PhyCNN $1.4386\times 10^{-3}$ $1.2763\times 10^{-3}$ 0.28% 0.33%
Table 6: Global error metrics for the six-story RC building case.
5.1.3 Peak response error analysis
The difference between the models becomes significantly more pronounced for peak response prediction. PhyULSTM achieves a mean peak error of $11.88\%$ and a median error of $9.78\%$, whereas PhyCNN exhibits substantially larger errors of $27.65\%$ and $32.76\%$, respectively. This shows that PhyCNN produces larger peak errors and a wider spread, while PhyULSTM maintains a more compact distribution. This result is significant because structural reliability and fragility assessments depend primarily on accurate estimation of extreme responses rather than average behavior.
@@ 6 Conclusion
This study proposes a novel Physics-Informed U-Net–LSTM (PhyULSTM) framework for the rapid prediction of seismic responses in nonlinear structural systems. This framework effectively addresses key limitations of standalone CNN and LSTM architectures in modeling nonlinear structural dynamics. The proposed framework is benchmarked against established physics-informed architectures, including PhyCNN and PhyLSTM, through two numerical case studies incorporating nonlinear stiffness behavior and a Bouc–Wen hysteretic model, as well as experimental validation using recorded sensor data from a mid-rise reinforced concrete building in San Bernardino. In all scenarios, the PhyULSTM demonstrated superior performance based on correlation coefficients, global error metrics, and peak response prediction accuracy. For cases where full-state measurements were available, 96.7% of predicted displacement responses achieved correlation coefficients exceeding 0.9, while in the acceleration-only training scenario, 94% exceeded this threshold. The mean and median peak response prediction errors were 8.92% and 5.76%, respectively, for the full-state case, and further reduced to 4.37% and 2.17% for the acceleration-only configuration. For the Bouc–Wen nonlinear hysteretic model, displacement prediction correlations ranged from 0.911 to 1.00, while velocity, acceleration, and restoring force predictions consistently closed to 0.99. For the mid-rise reinforced concrete building, the normalized prediction errors are close to zero: 99.89% of the third-floor predictions and 98.05% of the roof predictions fall within a ±5% error bound.
Beyond conventional statistical evaluation, the framework is further assessed for its ability to reproduce nonlinear hysteretic behavior, and for its ability to preserve critical frequency-domain characteristics essential for structural safety and reliability analysis. Through hysteresis loop reconstruction, Fourier amplitude spectrum analysis, and peak response error evaluation, and it is found that the model accurately captures nonlinear response evolution, dominant frequency components, and extreme structural demands. Importantly, the proposed architecture exhibits strong predictive capability under data scarce scenarios. For example, without explicit knowledge of structural parameters such as mass, stiffness, or damping matrices, the framework maintained reliable performance based only on acceleration measurements. This flexibility highlights its robustness for practical deployment in real-time structural health monitoring applications.
Overall, this study establishes PhyULSTM as a reliable, interpretable, and physically grounded metamodeling tool that advances the state of surrogate modeling for structural dynamics and provides a foundation for future developments in hybrid physics-informed machine learning. Future work will focus on enhancing model generalizability across various structural typologies, extending the framework toward inverse problems and parameter identification for structural health monitoring, and improving robustness under broader classes of seismic excitations.
Code Availability
The code corresponding to this work will be made publicly available after publication at the following repository: https://github.com/SuirthaBiswas/PhyULSTM